OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find g(4).

OpenStudy (broskishelleh):

@Twix47 Is this it?

OpenStudy (anonymous):

Same process as last time. Can you give it a try?

OpenStudy (anonymous):

so answer this time would be +4?

OpenStudy (anonymous):

First thing, can you write it out for me, how you would represent g(4)?

OpenStudy (anonymous):

If g(x) = x^2-4x g(4) = ...?

OpenStudy (anonymous):

still confused. every answer i put in it says im wrong so...

OpenStudy (anonymous):

\[g(x) = x^2-4x ~~~ g(4) = 4^2-4(4) =?\]

OpenStudy (skullpatrol):